# Privacy-Preserving Federated Learning with Differentially Private Hyperdimensional Computing

### v2

@@ 3.2 Differential Privacy
A random algorithm $M$ is said to satisfy $(\epsilon,\delta)$-DP if, for any two datasets $I_{1}$ and $I_{2}$ that differ by only a single data point, the following inequality holds:
@@
This condition ensures that the presence or absence of a single data point in the dataset has a limited impact on the outcome of the algorithm, thereby protecting individual privacy.
@@
The Gaussian Mechanism for any function $f:I\to\mathbb{R}^{D}$, which processes or transforms data, involves adding Gaussian noise with a mean of 0 and a variance of $\sigma_{dp}^{2}$ to the function’s output. This noise serves to obscure the contribution of any single data point, thereby ensuring DP.
@@
The sensitivity of a function $f$, denoted as $\Delta f$, is defined as the maximum difference in the output of $f$ when applied to any two datasets, $I_{1}$ and $I_{2}$, that differ by exactly one element. It is given by:
@@
This metric quantifies the extent to which a single data point can affect the outcome of the function.
@@
To ensure that a mechanism satisfies $(\epsilon,\delta)$-DP, the variance of the added noise, $\sigma_{dp}^{2}$, must meet the following condition:
@@
5: Main FedHD:
@@ 4 Research Methodology
In this study, we introduce the FedHDPrivacy framework, an innovative approach within the IoT domain that integrates FL, HD, and DP to securely facilitate the exchange of confidential information among clients. This framework utilizes HD as an explainable model to extract insights from training data and predict the labels of new samples. FL is employed to distribute the learning process across multiple clients, with each client training an HD model and then transmitting it to a central server. The server aggregates these local models to create a global model, which benefits from the combined knowledge of all clients, enhancing the accuracy of predictions for new data samples.
However, the exchange of information between clients and the server is susceptible to several types of attackers who can gain access to the ML models, including through eavesdropping, untrusted servers, server compromises, and malicious participants. Once these attackers have obtained the models, they can employ model inversion attacks to reconstruct training data or use membership inference attacks to determine whether a particular data sample was included in the training set. To counter these threats, DP is incorporated into the local models. We demonstrate that the noise added to each local model is sufficient to ensure the privacy of individual training samples. Furthermore, the global model, being an aggregation of noisy local models, does not expose any raw data, ensuring that server compromises cannot infer information about individual training samples. This approach also ensures that an untrusted server does not reveal client data, eavesdropping fails to extract confidential information from communication channels, and malicious participants only receive secured updates from the server.
In this section, we first outline an FL structure utilizing HD models to distribute the learning process among clients. Next, we introduce noise into each local model to enhance security. Thanks to the explainability of HD, the exact amount of noise required in each round can be precisely determined. Additionally, we compute the cumulative noise added across previous rounds and adjust the noise in the current round by accounting for the difference between the required noise and the noise already applied. This approach reduces the overall noise introduced into the global model during training, thereby improving its performance. In other words, our proposed method enables precise mathematical adjustments of noise levels, which is crucial for protecting privacy while maintaining model accuracy.
@@ 4.2 FedHDPrivacy
In the FL structure, class hypervectors are exchanged between clients and a central server. Without a privacy-preserving mechanism, this process can expose sensitive training data to model inversion and membership inference attacks. To protect this confidential information, Gaussian noise is added to the HD models. The variance of this noise is determined by Equation (8), which depends on several key parameters. The privacy loss threshold ($\delta$) must be less than the inverse of the number of training samples used in the model’s training. Following the approach by Saifullah et al. [57], we set $\delta$ as the inverse of the number of training samples. The sensitivity of the model ($\Delta f$) is specific to each application, and the privacy budget ($\epsilon$) is selected based on the desired privacy level.
Next, we discuss the noise required to secure both the local model ($\vec{\xi}_{k}^{r}$) for each client $k$, and the global model ($\vec{\xi}_{g}^{r}$), in each round $r$. Additionally, we account for the cumulative noise added in previous rounds, denoted as $\vec{\Psi}_{k}^{r-1}$ for the local models and $\vec{\Psi}_{g}^{r-1}$ for the global model. Finally, we compute the adjusted noise that needs to be added to maintain privacy. For the local models, this is calculated as $\vec{\Gamma}_{k}^{r}=\vec{\xi}_{k}^{r}-\vec{\Psi}_{k}^{r-1}$, and for the global model, it is $\vec{\Gamma}_{g}^{r}=\vec{\xi}_{g}^{r}-\vec{\Psi}_{g}^{r-1}$. This ensures that the privacy of all training samples is protected throughout the learning process. Finally, noise is added to both the local and global models, as demonstrated in Algorithm 2, to secure the FL structure and counteract model inversion and membership inference attacks.
@@ 5 Experimental Design
In industrial IoT systems, vast amounts of data are generated and require advanced analytical tools such as AI and ML to drive efficiency and improve production quality within smart manufacturing environments [58]. The geometric characteristics of manufactured parts, particularly their precision, play a crucial role in determining mechanical properties like stiffness, strength, and failure resistance [59]. Research has shown that even minor deviations from intended dimensions can result in defects, potentially leading to part failure [60]. Thus, accurate measurement, prediction, and control over these geometric dimensions are essential to maintain high product quality and prevent production defects. By integrating AI and ML into manufacturing workflows, companies can enhance their ability to predict, detect, and reduce potential defects, ensuring greater reliability and consistency in their products [61].
However, traditional centralized ML approaches that rely on transferring raw data to a central server raise critical data privacy and security concerns, especially in distributed IoT environments where sensitive manufacturing data may be compromised during transmission [48]. FL offers a promising alternative, enabling decentralized data processing that respects privacy needs in smart manufacturing settings. In FL systems, individual IoT devices or sensors train models locally on their data, transmitting only model parameters—rather than raw data—to a central server. These parameters are then aggregated into a global model, leveraging the distributed data while protecting privacy. This decentralized approach not only preserves data security but also harnesses the diversity of data across multiple devices, enhancing model robustness. For applications like quality control, where data labeling can be both costly and time-intensive [62], FL supports collaborative model building across manufacturing units, enabling more efficient and accurate predictions. This setup fosters continuous learning and monitoring, improving model performance while safeguarding sensitive information.
@@
In our case study, we implemented FL for quality control in smart manufacturing, focusing on 8 clients, as shown in Figure 4. Each client used a LASERTEC 65 DED hybrid CNC DMG machine to perform drilling tasks, producing 18 parts from 1040 steel blocks with dimensions of 76.2 mm $\times$ 76.2 mm $\times$ 76.2 mm. To evaluate part quality, the diameter of each drilled hole was measured, and Z-scores were calculated based on these measurements. These Z-scores categorized the parts into three groups: Nominal Drilling for Z-scores between -1 and 1, Under Drilling for Z-scores below -1, and Over Drilling for Z-scores above 1. Real-time data was collected using a Siemens Simatic IPC227E device connected to the DMG machine. This device recorded 15 critical process signals—including Load, Current, Torque, Commanded Speed, Control Differential (Diff), Power, Contour Deviation, Encoder Positions, Velocity Feed Forward, and Torque Feed Forward—at a frequency of 500 Hz during the manufacturing process. These signals, captured across five axes and the spindle, totaled 90 process signals per part.
@@
To support continuous FL operation, data from each part was segmented into samples using a windowing technique with a window size of 10, resulting in samples with 90 signals, each of length 10. Labels derived from Z-scores were applied to each sample, based on the drilled hole measurements of the part it came from. For each communication round, clients selected a balanced set of 500 samples, ensuring equal representation from each label category (Nominal Drilling, Under Drilling, and Over Drilling), and used these samples to update their local models. This process is repeated over 10 rounds to enhance model learning. The test dataset was constructed similarly by selecting balanced samples across all label categories from each client. The final global model’s performance was evaluated on the aggregated test data, ensuring robust and accurate results across the entire dataset.
@@ 6 Experimental Results
Figure 5 illustrates the relationship between distance and similarity among clients’ data points. The main diagonal of the figure presents the similarity between encoded samples as a function of the distance between them for each individual client. This provides valuable insight into how well the encoding preserves the internal structure of the data for each client. A high similarity for samples with shorter distances suggests that the encoding method is successfully capturing the inherent relationships within the client’s data. Notably, the data points for each client align on a line, indicating that the encoding strategy maps the data into the hyperspace effectively, without any outliers. For certain clients, such as Client 1, the similarity values range from 0 to 1, indicating a wide diversity between samples. In contrast, for Client 5, the similarity ranges from 0 to approximately 0.4, while the distances vary only between 8 and 16. This pattern suggests that Client 5’s samples are quite diverse, with little overlap in similarity, which highlights the distinctiveness of the data within this client’s set.
The lower triangle of Figure 5 shows the distribution of similarity between the hypervectors (encoded samples) for each pair of clients. This comparison highlights the level of similarity in the data representations across different clients. A higher similarity between clients may indicate overlap in the data characteristics, which could be relevant for tasks such as global model aggregation in FL. For instance, Clients 5 and 7 exhibit greater similarity, as their distribution peaks around a similarity value of 0.8, despite not having high similarity between the hypervectors of individual samples. In contrast, Clients 4 and 3 appear to have less similarity between their data, even though their individual samples, as shown in the main diagonal, exhibit similar internal patterns. This difference in cross-client similarity highlights the variation in data structure between different clients.
In contrast, the upper triangle of Figure 5 presents the distribution of distances between the actual samples from each pair of clients. This provides insight into the variance or diversity in the raw data across clients. Large distances between samples suggest significant differences in their data distributions, which may affect the model’s ability to generalize effectively across all clients. This combination of encoded sample similarities and raw data distances offers a holistic view of both the encoded representations and the underlying diversity in the system. For example, Clients 1 and 8 exhibit high distance values, with a distribution ranging from 12 to 16, indicating that the raw data for these clients is notably different. On the other hand, Clients 3 and 4 have more similar raw data points, as evidenced by a distance distribution peaking around 8.
@@
Figure 7: Cumulative and Required Noise Levels for Server Model. FedHDPrivacy framework noise levels for server models with 10, 50, and 100 clients, illustrating cumulative versus required noise. The framework controls noise per communication round, ensuring privacy without needing server-side noise. As the client count rises, cumulative noise variance decreases, balancing privacy with accuracy across rounds.
@@
Figure 10: Accuracy Comparison of FedHDPrivacy and Benchmark Federated Learning Frameworks. FedHDPrivacy effectively controls noise to prevent accuracy degradation across communication rounds, enabling it to outperform other federated learning frameworks, including FedAvg, FedNova, FedProx, FedAdam, and FedSGD. This balance of privacy and accuracy allows FedHDPrivacy to achieve higher performance in continuous learning scenarios.

### v3

@@ 3.2 Differential Privacy
A random algorithm $M$ satisfies $(\epsilon,\delta)$-DP if, for any two datasets $I_{1}$ and $I_{2}$ differing by only a single data point, the following condition holds:
@@
This ensures that the presence or absence of a single data point in the dataset has a minimal effect on the algorithm’s output, thereby safeguarding individual privacy.
@@
The Gaussian Mechanism applies to any function $f:I\to\mathbb{R}^{D}$ by adding Gaussian noise with mean 0 and variance $\sigma_{dp}^{2}$ to its output. This noise obscures the influence of any single data point, ensuring compliance with DP:
@@
The sensitivity of a function $f$, denoted as $\Delta f$, is the maximum difference in $f$’s output when applied to any two datasets $I_{1}$ and $I_{2}$ that differ by a single element:
@@
This metric quantifies the impact a single data point can have on the function’s output.
@@
For a mechanism to satisfy $(\epsilon,\delta)$-DP, the variance of the added Gaussian noise, $\sigma_{dp}^{2}$, must satisfy:
@@
These formulations provide the theoretical foundation for the DP mechanisms used in this work. Specifically, the Gaussian Mechanism and sensitivity calculations allow for precise control of noise introduction, ensuring privacy while minimizing its impact on model performance. By adhering to these principles, the proposed framework effectively balances privacy and utility, addressing the unique challenges of continuous learning in IoT-driven FL systems.
@@
5: Main Federated Hyperdimensional Computing:
@@ 4 Research Methodology
In this study, we propose the FedHDPrivacy framework, which integrates FL, HD, and DP to securely facilitate collaborative learning in IoT environments. The framework leverages HD as an explainable model to analyze training data and classify new samples. The learning process is distributed across multiple clients, where each client trains an HD model locally and sends its updates to a central server. The server aggregates these local models to form a global model, which combines the knowledge of all clients and improves predictive accuracy.
To address privacy vulnerabilities inherent in FL, such as model inversion and membership inference attacks, DP is incorporated into the framework. Specifically, noise is added to local models before transmission, ensuring that individual training samples cannot be inferred from the updates. The global model, as an aggregation of these noisy local models, does not expose sensitive information, even in the event of server compromise or malicious participant activity. This strategy also protects against eavesdropping during communication and mitigates risks associated with untrusted servers.
The methodology consists of two main components. First, HD models are used to distribute the learning process across clients, with local training updates aggregated to build the global model. Second, an adaptive DP mechanism is employed to manage noise levels. The explainability of HD allows for precise determination of the noise required in each communication round. By calculating the cumulative noise introduced in previous rounds, the framework adjusts the noise added in subsequent rounds to meet privacy guarantees while minimizing the impact on model performance. This approach ensures that the global model achieves a balance between privacy protection and accuracy, avoiding unnecessary degradation caused by excessive noise.
@@ 4.2 FedHDPrivacy
The FedHDPrivacy framework ensures privacy in FL by incorporating DP into the exchange of class hypervectors between clients and the central server. Without a privacy-preserving mechanism, this shared information is vulnerable to model inversion and membership inference attacks, which could reveal sensitive training data. To mitigate these risks, Gaussian noise is added to the HD models, with the variance of the noise calculated based on key privacy parameters: the privacy budget ($\epsilon$), the privacy loss threshold ($\delta$), and the sensitivity of the function ($\Delta f$). As described in Section 3.2, the noise variance is determined using Equation (8). The privacy loss threshold ($\delta$) is set as the inverse of the number of training samples, following the approach by Saifullah et al. [62], while the sensitivity ($\Delta f$) is application-specific, and the privacy budget ($\epsilon$) is chosen based on the desired level of privacy.
Next, we address the noise required to secure both the local model ($\vec{\xi}_{k}^{r}$) for each client $k$, and the global model ($\vec{\xi}_{g}^{r}$), in each round $r$. Additionally, we account for the cumulative noise introduced in previous rounds, denoted as $\vec{\Psi}_{k}^{r-1}$ for the local models and $\vec{\Psi}_{g}^{r-1}$ for the global model. Finally, we compute the adjusted noise that needs to be added to maintain privacy. For the local models, this is calculated as $\vec{\Gamma}_{k}^{r}=\vec{\xi}_{k}^{r}-\vec{\Psi}_{k}^{r-1}$, and for the global model, it is $\vec{\Gamma}_{g}^{r}=\vec{\xi}_{g}^{r}-\vec{\Psi}_{g}^{r-1}$. This ensures that the privacy of all training samples is protected throughout the learning process. Additionally, noise is added to both the local and global models, as demonstrated in Algorithm 2, to secure the FL structure and counteract model inversion and membership inference attacks.
In the first round of FL, clients independently train their local HD models using $L$ training samples and share their noisy class hypervectors with the server. Theorem 2 defines the amount of noise required to ensure privacy while maintaining accuracy, with Proof 1 deriving the sensitivity and variance of the noise.
@@
Once the local models are aggregated at the server, Theorem 3 establishes that no additional noise is required for the global model in the first round. This conclusion is demonstrated in Proof 2, which is further supported by Lemma 1 and Proof 3.
@@
In continuous FL scenarios, new data is introduced in each round, necessitating noise adjustments to ensure cumulative privacy guarantees. For round $r\geq 2$, the noise required for local models is defined in Theorem 4, which considers both the cumulative noise from previous rounds and the additional noise needed for the current round, as demonstrated in Proof 4.
@@
After aggregating the local models in round $r\geq 2$, Theorem 5 establishes that no additional noise is required for the global model, as the privacy guarantees are maintained by the noise introduced in the local models. This result is rigorously demonstrated in Proof 5.
@@
Lemma 2 ensures that the cumulative noise from clients is sufficient to protect the global model across rounds, eliminating the need for additional noise from the server. This result is further supported and demonstrated in Proof 6.
@@ 5 Experimental Design
In industrial IoT systems, vast amounts of data are generated and require advanced analytical tools such as AI and ML to drive efficiency and improve production quality within smart manufacturing environments [63]. The geometric characteristics of manufactured parts, particularly their precision, play a crucial role in determining mechanical properties like stiffness, strength, and failure resistance [64]. Research has shown that even minor deviations from intended dimensions can result in defects, potentially leading to part failure [65]. Thus, accurate measurement, prediction, and control over these geometric dimensions are essential to maintain high product quality and prevent production defects [66].
5.1 Datasets and Data Preprocessing
@@
In our case study, we implemented FL for quality control in smart manufacturing, focusing on 8 clients, as shown in Figure 4. Each client used a LASERTEC 65 DED hybrid CNC DMG machine to perform drilling tasks, producing 18 parts from 1040 steel blocks with dimensions of 76.2 mm $\times$ 76.2 mm $\times$ 76.2 mm. The diameter of each drilled hole was measured, and Z-scores were calculated to categorize parts into three groups: Nominal Drilling ($-1\leq Z\leq 1$), Under Drilling ($Z<-1$), and Over Drilling ($Z>1$).
Real-time data was collected using a Siemens Simatic IPC227E device connected to the DMG machine. This device recorded 15 critical process signals—including Load, Current, Torque, Commanded Speed, Control Differential (Diff), Power, Contour Deviation, Encoder Positions, Velocity Feed Forward (FF), and Torque FF—at a frequency of 500 Hz during the manufacturing process. These signals, captured across five axes and the spindle, totaled 90 process signals per part.
Table 2: Summary of transformer and convolutional neural network architectures
Model Details
Transformer Input Projection: Linear layer (10 → 128)
Positional Encoding: Learnable embeddings
2 Transformer Encoder layers: 4 heads, FFN size 256, ReLU, Dropout=0.1
2 Fully Connected layers: 128 → 8
Convolutional Neural Network 2 Convolutional layers: 32 and 64 filters (3x3 kernels, ReLU activation)
MaxPooling layer (2x2, stride 2)
3 Fully Connected layers: 14400 → 1024 → 1024 → 8
To generate samples, each part’s 90 recorded signals were segmented into subsignals using a sliding window approach with a window size of 10. This resulted in multiple samples per part, where each sample contained 90 signals of length 10. The label assigned to each sample was inherited from the Z-score category of the part from which the data was recorded.
5.2 Experimental Setup
@@
For balanced evaluation, we ensured equal representation from each label category (Nominal Drilling, Under Drilling, and Over Drilling) by selecting an equal number of samples per client. The dataset was split such that 80% of samples were used for training and 20% for testing the global model. During FL training, each client’s training set was divided into 10 subsets. In each communication round, a single subset was used to update the local model, ensuring 10 communication rounds in total. The test dataset was constructed by combining all clients’ test samples. The global model’s performance was evaluated on the aggregated test data, ensuring robust and accurate results.
5.3 Hyperparameters and Evaluation Metrics
We conducted a sensitivity analysis to evaluate the proposed method under different settings by tuning key hyperparameters. Noise variance was adjusted based on the privacy budget ($\epsilon$) and tracked across rounds to minimize cumulative effects. The hypervector size and privacy budget ($\epsilon$) were varied to examine their influence on model accuracy, training time, and energy consumption. The number of rounds was fixed at 10 to balance communication overhead and model convergence. To assess model performance, we used accuracy, which measures the proportion of correctly classified samples in the test set; training time, which represents the total duration required for model convergence; and energy consumption, which was measured at the device level during training to evaluate efficiency.
5.4 Comparison with Benchmark Frameworks
We compared FedHDPrivacy against FedAvg, FedNova, FedProx, and FedOpt. These frameworks were trained using two ML architectures: a Transformer and a Convolutional Neural Network (CNN), detailed in Table 2. Each benchmark was tested under identical conditions using the same dataset and experimental setup. The performance of these frameworks was analyzed across the selected metrics to highlight the advantages of the proposed method in terms of privacy preservation, accuracy, and efficiency.
@@ 6 Experimental Results
6.1 Sensitivity Analysis
Figure 5 illustrates the relationship between distance and similarity among clients’ data points. The main diagonal of the figure presents the similarity between encoded samples as a function of the distance between them for each individual client. This provides valuable insight into how well the encoding represents the internal structure of the data for each client. A high similarity for samples with shorter distances suggests that the encoding method is successfully capturing the inherent relationships within the client’s data. Notably, the data points for each client align on a line, indicating that the encoding strategy maps the data into the hyperspace effectively, without any outliers. For certain clients, such as Client 1, the similarity values range from 0 to 1, indicating a wide diversity between samples. In contrast, for Client 5, the similarity ranges from 0 to approximately 0.4, while the distances vary only between 8 and 16. This pattern suggests that Client 5’s samples are quite diverse, with little overlap in similarity, which highlights the distinctiveness of the data within this client’s set.
The lower triangle of Figure 5 shows the distribution of similarity between the hypervectors (encoded samples) for each pair of clients. This comparison highlights the level of similarity in the data representations across different clients. A higher similarity between clients may indicate commonality in the data characteristics, which could be relevant for tasks such as global model aggregation in FL. For instance, Clients 5 and 7 exhibit greater similarity, as their distribution peaks around a similarity value of 0.8, despite not having high similarity between the hypervectors of individual samples. In contrast, Clients 4 and 3 appear to have less similarity between their data, even though their individual samples, as shown in the main diagonal, exhibit similar internal patterns. This difference in cross-client similarity highlights the variation in data structure between different clients.
In contrast, the upper triangle of Figure 5 presents the distribution of distances between the actual samples from each pair of clients. This provides insight into the divergence or diversity in the raw data across clients. Large distances between samples suggest significant differences in their data distributions, which may affect the model’s ability to generalize effectively across all clients. This combination of encoded sample similarities and raw data distances offers a holistic view of both the encoded representations and the underlying diversity in the system. For example, Clients 1 and 8 exhibit high distance values, with a distribution ranging from 12 to 16, indicating that the raw data for these clients is notably different. On the other hand, Clients 3 and 4 have more similar raw data points, as evidenced by a distance distribution peaking around 8.
@@ 6.1 Sensitivity Analysis
Figure 7: Cumulative and required noise levels for the server’s model. FedHDPrivacy framework noise levels for server models with 10, 50, and 100 clients, illustrating cumulative versus required noise. The framework controls noise per communication round, ensuring privacy without needing server-side noise. As the client count rises, cumulative noise variance decreases, balancing privacy with accuracy across rounds.
@@
Figure 10: Accuracy of FedHDPrivacy as a function of hypervector size and privacy budget. Larger hypervector sizes improve accuracy by increasing the model’s capacity, while lower privacy budgets ($\epsilon$), indicating higher privacy levels, reduce accuracy.
Figure 11: Energy efficiency and training time as a function of hypervector size. Larger hypervector sizes require more energy and time for the training process.
@@
For the given task with a fixed number of clients and training samples, the main hyperparameters of the proposed method are hypervector size and privacy budget. Figure 10 illustrates the accuracy of FedHDPrivacy as a function of hypervector size and privacy budget. Larger hypervector sizes improve accuracy by increasing the model’s capacity, while lower privacy budgets ($\epsilon$), which indicate higher privacy levels, reduce accuracy. When the hypervector size is 1K, the accuracy is below 54%, even with low noise levels, demonstrating that the model requires a larger hyperspace dimension for effective training. By increasing the hypervector size, accuracy improves significantly. For example, when the hypervector size is 10K, the accuracy exceeds 63%, even at high privacy levels. Furthermore, decreasing the privacy level (i.e., increasing $\epsilon$) reduces the noise, thereby increasing the model’s accuracy. For instance, when $\epsilon=10$, the accuracy surpasses 70%. Finally, at a hypervector size of 60K, the accuracy improves further. In this case, by applying the minimum amount of noise, the accuracy reaches 74.21%. Thus, increasing the hypervector size enhances accuracy.
However, increasing the hypervector size also increases training time and energy consumption, as shown in Figure 11. For example, when the hypervector size is 1K, the training time and energy consumption are 32.56 seconds and 1.5 KJ, respectively. By increasing the hypervector size to 30K, the training time and energy consumption rise to 82.88 seconds and 13.98 KJ, respectively. This indicates that training time increases by 154.55% and energy consumption increases by 832% as the hypervector size grows. To balance accuracy, privacy, and efficiency, we selected a hypervector size of 10K and set $\epsilon=10$ for benchmarking against other FL frameworks.
Figure 12: Accuracy comparison of FedHDPrivacy and benchmark FL frameworks using (a) transformer and (b) convolutional neural network over communication rounds. FedHDPrivacy outperforms FedAvg, FedNova, FedProx, and FedOpt by effectively managing noise to maintain accuracy across communication rounds, demonstrating superior performance in continuous learning scenarios.
Table 3: Comparison of false positive and false negative rates across different models.
Federated Learning Framework Machine Learning Model False Positive Rate $\downarrow$ False Negative Rate $\downarrow$
FedHDPrivacy Hyperdimensional Computing 14.63% 29.27%
FedAvg [8] Transformer 26.68% 53.36%
FedNova [32] Transformer 27.67% 55.34%
FedProx [31] Transformer 31.85% 63.69%
FedOpt [33] Transformer 33.33% 66.67%
FedAvg [8] Convolutional Neural Network 24.62% 49.23%
FedNova [32] Convolutional Neural Network 22.61% 45.22%
FedProx [31] Convolutional Neural Network 25.33% 50.66%
FedOpt [33] Convolutional Neural Network 33.33% 66.67%
6.2 Benchmark Comparison